OpenStudy (kpop4life123):

An average clementine is shaped like a sphere and has a diameter of 3.5 inches. What is the volume of the fruit? Use 3.14 for pi. Round your answer to the nearest hundredth.

OpenStudy (kpop4life123):

Can u help me on this too @Chealyn98

OpenStudy (kpop4life123):

20.45 cubic inches 21.34 cubic inches 22.44 cubic inches 23.78 cubic inches

OpenStudy (kpop4life123):

??????

OpenStudy (kittiwitti1):

Why the question marks?

OpenStudy (kpop4life123):

u wrote something then u deleted it

OpenStudy (kittiwitti1):

\[V=4\pi r^{3},\text{ }r=\frac{d}{2}=\frac{3.5}{2}\approx1.75\rightarrow V=4\pi(1.75)^{3}=?\]Really? Maybe your computer is glitching lol

OpenStudy (kpop4life123):

lol but i just bought it

OpenStudy (kittiwitti1):

Your internet then >_> Have you got the answer? :p

OpenStudy (kpop4life123):

no:/ im really bad at math

OpenStudy (kittiwitti1):

Simplify the equation at the end of what I gave you:\[V=4\pi(1.75)^{3}=?\]Hint: use a calculator

OpenStudy (kittiwitti1):

Well it says use 3.14 for pi so do this:\[V=4\times3.14(1.75)^{3}=?\]

OpenStudy (chealyn98):

sleepyjess Mathlete 2 d=3.5 r=d/2 r=3.5/2

OpenStudy (kpop4life123):

664.697056333927395 i think i wrote it on the cal. right

OpenStudy (kittiwitti1):

Did you use the equation I gave you?

OpenStudy (kittiwitti1):

the last one btw

OpenStudy (kpop4life123):

let me check again

OpenStudy (kpop4life123):

67.31375

OpenStudy (kpop4life123):

@kittiwitti1

OpenStudy (chealyn98):

67.31375 is the answer

OpenStudy (kittiwitti1):

Yes! :D

OpenStudy (chealyn98):

Your welcome

OpenStudy (kpop4life123):

but its not one of the options?????????

OpenStudy (kpop4life123):

20.45 cubic inches 21.34 cubic inches 22.44 cubic inches 23.78 cubic inches

OpenStudy (kittiwitti1):

hmmm

OpenStudy (kittiwitti1):

that's weird...

OpenStudy (kittiwitti1):

http://openstudy.com/study#/updates/544914ebe4b0d6442915267a sorry it was squared not cubed! idk google told me cubed >_<